# Character theoretic techniques for nonabelian partial difference sets

### v1

@@ 3.2. Results that hold when a prime $p$ divides $v$ but not $\sqrt{\Delta}$
We now prove $\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}\leqslant{\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})$. We must show first that there exists a set of basis vectors $v_{1},\dots,v_{k}$ of $M_{\mathfrak{F}}(\theta_{\alpha})$ which do not vanish in $\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$. For this, we use the Elementary Divisor Theorem (see [22, Chapter III, Theorem 7.8]), which states that there exists set of vectors $v_{1},\dots,v_{n}$ of $V_{\mathfrak{F}}(\theta_{\alpha})$ so that for some $r_{1},\dots,r_{k}\in\mathfrak{R}$, the vectors $r_{1}v_{1},\dots,r_{k}v_{k}$ form a basis for $M_{\mathfrak{F}}(\theta_{\alpha})$. Now, if $r_{i}v_{i}$ is an eigenvector of $\mathfrak{F}(D)$, then $v_{i}$ itself must be an eigenvector of $\mathfrak{F}(D)$, for $r_{i}$ is a scalar. Therefore, $v_{i}$ is in the span of the basis, so we may choose $r_{i}$ to be 111. Therefore, $v_{1},\dots,v_{k}$ is a basis for $M_{\mathfrak{F}}(\theta_{\alpha})$. Since $\mathfrak{p}M_{\mathfrak{F}}$ is defined as the set of $\mathfrak{p}$ multiples of the basis vectors of $M_{\mathfrak{F}}$, then $\overline{v_{i}}\neq 0$ for each $v_{i}$. Each $\overline{w}\in\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$ may be expressed as a sum of the $\overline{v_{i}}$ for we may take some pre-image $w$ of $\overline{w}$, and then $a_{1}v_{1}+\dots+a_{k}v_{k}=w$, so $\overline{a_{1}}\overline{v_{1}}+\dots+\overline{a_{k}}\overline{v_{k}}=\overline{w}$ as needed.
@@
Let $\xi,\chi$ be linear characters of order $p^{a},q^{b}$ respectively such that $p$ and $q$ are coprime to $\sqrt{\Delta}$. Then, $\xi(D)=\chi(D)$. In addition, all linear characters of order coprime to $\sqrt{\Delta}$ are equivalent over $D$.
@@
In order to prove Corollary 3.10, which is useful when ruling out the possibility of certain groups containing a PDS, we introduce the set of primes $\Pi_{\alpha}$ and $\Pi_{\beta}$ with products $\pi_{\alpha}$ and $\pi_{\beta}$ defined so that $\pi_{\alpha}$ and $\pi_{\beta}$ are the largest integers dividing $v_{\alpha}$ and $v_{\beta}$, respectively, such that $\gcd(\sqrt{\Delta},\pi_{\alpha})=\gcd(\sqrt{\Delta},\pi_{\beta})=1$.
@@
Suppose that $\xi\in\mathcal{L}$ has order dividing $\pi_{\alpha}$. Then, each $\xi\in\mathcal{L}$ with order coprime to $\sqrt{\Delta}$ has an order dividing $\pi_{\alpha}$.
@@
For this next theorem, and for the remainder of the section, let $H$ be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$, and let $N$ be the intersection of the kernels of the characters in $H$. The following theorem is a consequence of Corollary 3.9. The reader may compare Theorem 3.11 to [25, Theorem 2.7].
@@
If $a\not\in N$, then for $\xi\in H$, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$.
@@
[…]
@@
For $a\not\in N$, the greatest common divisor of $|C_{G}(a)|$ and at least one of $\pi_{\alpha},\pi_{\beta}$ is 111.
@@
[…]
@@ 3.3. Intersections of $D$ with certain normal subgroups
Let $G^{\prime}=[G,G]$ be the derived subgroup of $G$, and let $\mathcal{L}$ be the group of linear characters of $G$; note that $\mathcal{L}\cong G/G^{\prime}$. Recall that we defined $H$ to be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$ and $N$ to be the intersection of the kernels of the characters in $H$. In the case when $\gcd(v,\sqrt{\Delta})=1$, we see that $|\mathcal{L}|$ is coprime to $\sqrt{\Delta}$, so we can choose $H$ to be $\mathcal{L}$, in which case $N=G^{\prime}$. We note first the following result.
@@
[…]
@@
If $N<G$, then
$|N\cap D|=\frac{k-\theta_{\alpha}}{|H|}+\theta_{\alpha}.$
@@
First, we apply Lemma 3.4(iii) to $N$ to obtain
@@
[…]
$|Na\cap D|=\frac{k-\theta_{\alpha}}{|H|}.$
@@
If $G$ contains a regular $(v,k,\lambda,\mu)$-PDS and $\gcd(v,\sqrt{\Delta})=1$, then
@@ 3.4. Example applications
Let $q$ be even, and assume that a group $G$ acts regularly on the point set of a GQ with order $(q,q^{2}-q)$. This implies the existence of a $((q+1)(q^{2}-q+1),q(q^{2}-q+1),q-1,q^{2}-q+1)$-PDS; in this case, the eigenvalues $\theta_{1},\theta_{2}$ are $q-1$, $-q^{2}+q-1$, and $\sqrt{\Delta}=q^{2}$. Note that if $p|\sqrt{\Delta}$, then $p\not|\;v=(q+1)(q^{2}-q+1)$, so we may apply Theorem 3.13. Thus, we let $p$ divide $v$. Now, $k-\theta_{1}=q^{3}-q^{2}+1$ and $k-\theta_{2}=(q+1)(q^{2}-q+1)$. We show that $v_{1}=k-\theta_{1}$ and $v_{2}=\frac{k-\theta_{2}}{\mu}$. So, suppose $p|\mu$ and $p|k-\theta_{1}$. Then mod $p$,
0 0 $\equiv q^{3}-q^{2}+1$
@@
So $p|q$ and $p|1-q$, forcing $p=1$, so $\mu$ cannot divide $k-\theta_{1}$. Therefore, $v_{1}=q^{3}-q^{2}+1$ and $v_{2}=q+1$.
[…]
@@
[…]
@@
[…]
Now, let $a\in G$ be an element of order divisible by 333 such that $a\not\in[G,G]$. Such an element must exist, for 333 divides $|\mathcal{L}|$. In addition, let $\xi\in\mathcal{L}$. Then, by Theorem 3.11, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$. By Lemma 3.2, $\xi(D)=\pm w$. Set $w=3m+1$ for some $m\in\mathbb{Z}$. Therefore,
@@
Since $a$ has order divisible by 333, then 333 divides $|C_{G}(a)|$, so $6m+2-(\pm 1)\equiv 0\pmod{3}$. Since $w\neq 0\pmod{3}$, then $2-(\pm 1)\equiv 0\pmod{3}$. It follows that $\xi(D)=-w$, so that 2(±1)2-($\pm$ 1)2 - ( ± 1 ) equals 333. ∎
@@ 4. Computational methods
These methods, combined with Theorem 3.11, makes it rather quick to search for groups which could contain a PDS just by utilizing the irreducible characters of the group. In some cases, simply knowing the values $|D\cap h^{G}|\pmod{p^{\ell}}$ can rule out the existence of a PDS for a certain parameter sets inside of a nonabelian group, such as in Example 4.2. In other cases, targeted information about $|D\cap h^{G}|$ can be used to quickly construct partial difference sets even inside of groups with up to 100010001000 elements, using the search method described in [2].
@@ 4.1. Modular computation of conjugacy class intersection sizes for PDSs
For $h\in G$ such that $h\neq 1$, suppose that $p^{\ell}$ does not divide $|C_{G}(h)|$. Then, over the ring $R$,
@@
There does not exist an invertible (112,75,50)(112,75,50)( 112 , 75 , 50 )-DS in a nonabelian group $G$ of order 112112112. We suppose for contradiction that there does exist such a DS in a nonabelian group $G$ of order 112112112. This parameter set has eigenvalues $\theta_{1}=5$ and $\theta_{2}=-5$. By Lemma 3.2, if $D$ exists then $\overline{\chi(D)}\equiv 0\pmod{5}$. Therefore, by Lemma 4.1, $\Phi(g)\equiv 75-5+\sum_{\chi\in{\rm Irr}(G)}(\chi(1)\theta_{2})\chi(g)\equiv 0\pmod{5}$. Finally, there is no centralizer $C_{G}(g)$ of order divisible by 555, since 555 does not divide 112112112. Therefore, for all $g\in G$, $|g^{G}\cap D|\equiv 0\pmod{5}$.
[…]
An analogous computation for each group verifies that there are no nonabelian groups of order 112112112 containing a regular (112,75,50)(112,75,50)( 112 , 75 , 50 )-DS. A direct verification may be done in GAP[15], using the following command:
@@
Let $D$ be a regular PDS. If $p^{\ell}$ divides both $\theta_{1}$ and $\theta_{2}$, then $k\equiv 0\pmod{p^{\ell}}$. In particular, $k\equiv 0\pmod{\gcd(\theta_{1},\theta_{2})}$.
@@
Let $D$ be a regular $PDS$ with eigenvalues $\theta_{1},\theta_{2}$ such that $\gcd(\theta_{1},\theta_{2})\neq 1$. If $p^{\ell}$ divides $\gcd(\theta_{1},\theta_{2})$ and $p^{\ell}$ does not divide $v$, then for each conjugacy class $h^{G}$,
@@
We explore the structure of a (183,112,66,72)(183,112,66,72)( 183 , 112 , 66 , 72 )-PDS $D$ in a group $G$ of order 183183183. The eigenvalues of this PDS are $\theta_{1}=4$ and $\theta_{2}=-10$. Then, $\gcd(4,-10)=2$, but 222 does not divide 183183183. Therefore, by Lemma 4.4, for every $h\in G$, we have $|h^{G}\cap D|\equiv 0\pmod{2}$. By a computation in GAP[15], there is one nonabelian group given by SmallGroup(183, 1) with two conjugacy classes $h_{1}^{G},h_{2}^{G}$ of order 616161, 202020 conjugacy classes of order 333, and one conjugacy class of order 111. If $x\in h_{1}^{G},h_{2}^{G}$, then $x\not\in G^{\prime}$ by a computation in GAP[15]. Moreover, every prime $p$ dividing 183183183 does not divide 141414, and $G$ must have a subgroup of linear characters, for it is solvable by Feit-Thompson. So, by Theorem 3.11, $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|\in\{\frac{112-4}{3},\frac{112+10}{3}\}$. Only $\frac{112-4}{3}=36$ is integral, so $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|=36$. Therefore, 11272=40112-72=40112 - 72 = 40 elements must be partitioned among 202020 conjugacy classes. The maximum size of each conjugacy class is 333, and each class must have $0\pmod{2}$ elements, so each of the remaining conjugacy classes contain 222 elements of $D$.
@@
Combining Lemma 4.1 and Theorem 3.11 lets us build a fast algorithm to solve for all values $|D\cap h^{G}|$ even in comparatively large groups of order up to 100010001000. We briefly describe the concept of such an algorithm, and then provide the pseudocode for an implementation.
@@
We show that there does not exist a (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 )-PDS inside of $G=$ SmallGroup(100, 3). By a computation in GAP[15], $G$ contains a conjugacy class $h^{G}$ of order 444 such that $g,g^{-1}\in h^{G}$ for all $g\in h^{G}$. Moreover, the centralizer $C_{G}(h)$ has order 252525. Since the eigenvalues corresponding to (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 ) are 888 and 3-3- 3, by Lemma 4.1,
@@
The only nonnegative integer less than or equal to 444 that is congruent to $1\pmod{11}$ is 111, so $|h^{G}\cap D|=1$. Then, by Lemma 4.6, $G$ cannot contain a (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 )-PDS.

### v2

@@ 3.2. Results that hold when a prime $p$ divides $v$ but not $\sqrt{\Delta}$
We now prove $\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}\leqslant{\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})$. We must show first that there exists a set of basis vectors $v_{1},\dots,v_{k}$ of $M_{\mathfrak{F}}(\theta_{\alpha})$ which do not vanish in $\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$. For this, we use the Elementary Divisor Theorem (see [22, Chapter III, Theorem 7.8]), which states that there exists set of vectors $v_{1},\dots,v_{n}$ of $V_{\mathfrak{F}}(\theta_{\alpha})$ so that for some $r_{1},\dots,r_{k}\in\mathfrak{R}$, the vectors $r_{1}v_{1},\dots,r_{k}v_{k}$ form a basis for $M_{\mathfrak{F}}(\theta_{\alpha})$. Now, if $r_{i}v_{i}$ is an eigenvector of $\mathfrak{F}(D)$, then $v_{i}$ itself must be an eigenvector of $\mathfrak{F}(D)$, for $r_{i}$ is a scalar. Therefore, $v_{i}$ is in the span of the basis, so we may choose $r_{i}$ to be $1$. Therefore, $v_{1},\dots,v_{k}$ is a basis for $M_{\mathfrak{F}}(\theta_{\alpha})$. Since $\mathfrak{p}M_{\mathfrak{F}}$ is defined as the set of $\mathfrak{p}$ multiples of the basis vectors of $M_{\mathfrak{F}}$, then $\overline{v_{i}}\neq 0$ for each $v_{i}$. Each $\overline{w}\in\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$ may be expressed as a sum of the $\overline{v_{i}}$ for we may take some pre-image $w$ of $\overline{w}$, and then $a_{1}v_{1}+\dots+a_{k}v_{k}=w$, so $\overline{a_{1}}\overline{v_{1}}+\dots+\overline{a_{k}}\overline{v_{k}}=\overline{w}$ as needed.
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$. Let $\xi,\chi$ be nontrivial linear characters of $G$ of prime-power orders $p^{a},q^{b}$, respectively, such that $p$ and $q$ are coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$. Then, $\xi(D)=\chi(D)$. Moreover, all linear characters of order coprime to $\sqrt{\Delta}$ are equivalent over $D$.
@@
Definition 3.10.
Recall from Lemma 3.6 that, if there exists a prime $p$ dividing $v$ but not $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, then $k-\theta_{1}$ and $k-\theta_{2}$ are coprime, and so there exists a factorization $\mu=\mu_{1}\mu_{2}$ such that $v=v_{1}v_{2}$, where $v_{\alpha}=(k-\theta_{\alpha})\mu_{\alpha}$ for $\alpha=1,2$, and $p$ divides only one of $v_{1}$ or $v_{2}$. We define the set of primes $\Pi_{\alpha}$ and $\Pi_{\beta}$ ($\{\alpha,\beta\}=\{1,2\}$) with products $\pi_{\alpha}$ and $\pi_{\beta}$ defined so that $\pi_{\alpha}$ and $\pi_{\beta}$ are the largest integers dividing $v_{\alpha}$ and $v_{\beta}$, respectively, such that $\gcd(\sqrt{\Delta},\pi_{\alpha})=\gcd(\sqrt{\Delta},\pi_{\beta})=1$.
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\mathcal{L}$ be the group of linear characters of $G$. Suppose $\pi_{\alpha}$ is defined as in Definition 3.10, and suppose that $\xi\in\mathcal{L}$ has order dividing $\pi_{\alpha}$. Then, each $\xi\in\mathcal{L}$ with order coprime to $\sqrt{\Delta}$ has an order dividing $\pi_{\alpha}$.
@@
The following theorem is a consequence of Corollary 3.9. The reader may compare Theorem 3.12 to [25, Theorem 2.7].
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$. If $a\not\in N$, then for nontrivial $\xi\in H$, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$.
@@
Let $\xi\in H$ be a linear character of order coprime to $|G|$. Let $\chi$ be an arbitrary irreducible character of $G$. We have
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, let $N$ be the intersection of the kernels of the characters in $H$, and let $\pi_{\alpha},\pi_{\beta}$ be defined as in Definition 3.10. For $a\not\in N$, the greatest common divisor of $|C_{G}(a)|$ and at least one of $\pi_{\alpha},\pi_{\beta}$ is $1$.
@@
Since $p$ is coprime to $\sqrt{\Delta}$, and $p$ divides $|\mathcal{L}|$, we have that $H$ is nontrivial. Thus, $N\neq G$, so the hypotheses for Lemma 3.13 apply. Suppose there exists $q$ dividing $|N_{G}(P)|$ such that $q\in\Pi_{\beta}$. Then, there is an element $g\in N_{G}(P)$ of order $q$. By [18, Theorem 3.5], $P=C_{P}(g)[\langle g\rangle,P]$. Each element in $C_{P}(g)$ centralizes $g$, so each element $a\in C_{P}(g)$ contains $g$ in its centralizer. Thus, for some $m$, $\gcd(|C_{G}(a)|,\pi_{\beta})=q\cdot m$, and since $a\in C_{P}(g)$ and $a\in C_{G}(a)$, then the order of $a$ divides both $|P|$ and $|C_{G}(a)|$. Thus, $\gcd(|C_{G}(a)|,|P|)>1$, forcing $\gcd(|C_{G}(a)|,\pi_{\alpha})>1$. Therefore, by the contrapositive of Lemma 3.13, $a$ is in $N$. Thus, $P=C_{P}(g)[\langle g\rangle,P]\subseteq NG^{\prime}\subseteq N$ (where $G^{\prime}=[G,G]$), given that $G^{\prime}\subseteq N$, since for each $\xi\in H$, we must have $G^{\prime}\subseteq\ker\xi$. So, $P\subseteq N$. But, $p$ divides $|\mathcal{L}|$, a contradiction. Therefore, $\gcd(|N_{G}(P)|,\pi_{\beta})=1$.
Now, suppose in addition that $G$ is a solvable group. By Hall’s Theorem, there is a subgroup $K$ of $G$ which is a $\pi_{\alpha}\pi_{\beta}$-group. Moreover, since $G$ is solvable, $K$ is solvable, so there exists a subgroup $U$ of $K$ which is a $\pi_{\alpha}$-group. Since $p\in\Pi_{\alpha}$, then $|P|$ divides $\pi_{\alpha}$. Therefore, $P$ is contained in some Hall subgroup, $L$. Any two Hall subgroups are conjugate, and thus there exists $g\in G$ such that $g^{-1}Lg=K$. So, $K$ contains a Sylow $p$-subgroup, $g^{-1}Pg$. Therefore, we are justified in assuming that $P\subseteq K$, forcing $P\subseteq U$. Thus by our preceding argument, $N_{K}(P)=N_{K}(U)$. Therefore, using Sylow III,
@@ 3.3. Intersections of $D$ with certain normal subgroups
Let $G^{\prime}=[G,G]$ be the derived subgroup of $G$, and let $\mathcal{L}$ be the group of linear characters of $G$; note that $\mathcal{L}\cong G/G^{\prime}$. We will define $H$ to be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$ and $N$ to be the intersection of the kernels of the characters in $H$. In the case when $\gcd(v,\sqrt{\Delta})=1$, we see that $|\mathcal{L}|$ is coprime to $\sqrt{\Delta}$, so we can choose $H$ to be $\mathcal{L}$, in which case $N=G^{\prime}$. We note first the following result.
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$. If $|H|\neq\{1\}$ and $\xi(D)=\theta_{\alpha}$ for some nonprincipal linear character $\xi$ in $H$, then $\xi(D)=\theta_{\alpha}$ for all nonprincipal $\xi\in H$. Moreover, for all $a\in G\backslash N$,
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$. If $N<G$, then
$|N\cap D|=\frac{k-\theta_{\alpha}}{|H|}+\theta_{\alpha},$
where $\xi(D)=\theta_{\alpha}$ for all nonprincipal $\xi\in H$.
@@
First, we apply Lemma 3.4(iii) to NN to obtain
@@
Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$. If $N<G$, and $a\in G\backslash N$, then
$|Na\cap D|=\frac{k-\theta_{\alpha}}{|H|},$
where $\xi(D)=\theta_{\alpha}$ for all nonprincipal $\xi\in H$.
@@
Assume the group $\mathcal{L}$ of linear characters of $G$ is nontrivial. If $G$ contains a regular $(v,k,\lambda,\mu)$-PDS, $\gcd(v,\sqrt{\Delta})=1$, and $\xi(D)=\theta_{\alpha}$ for a nontrivial $\xi\in\mathcal{L}$, then
@@ 3.4. Example applications
[…]
$0$ $\equiv q^{3}-q^{2}+1$
@@
So $p|q$ and p|1qp|1-q, forcing $p=1$, so $\mu$ cannot divide $k-\theta_{1}$. Therefore, $v_{1}=q^{3}-q^{2}+1$ and $v_{2}=q+1$.
Note that $v$ is odd, so if $G$ acts regularly on the points of this GQ, then $G$ is solvable by Feit-Thompson. Therefore, if there is some linear character of order $p$, then either $v_{1}\equiv 0$ and $v_{2}\equiv 1\pmod{p}$ or $v_{1}\equiv 1$ and $v_{2}\equiv 0\pmod{p}$ by Theorem 3.14. If $v_{1}\equiv 0$ and $v_{2}\equiv 1$, then $q\equiv 0$, forcing $v_{1}\equiv 1$, and thus $p=1$. If $v_{1}\equiv 1$ and $v_{2}\equiv 0$, then $q\equiv-1$, so $v_{1}\equiv(-1)^{3}-(-1)^{2}+1\equiv-1$, forcing $1\equiv 1$ and thus $p=2$. Therefore, $2|q+1$, and so $q$ is odd, a contradiction. Therefore, no such $G$ exists. ∎
@@
If there exists a $(4w^{2}-1,2w^{2},w^{2},w^{2})$-PDS $D$ in a group $G$, then $w\equiv 1\pmod{3}$, all linear characters of $G$ evaluate to w-w on $D$, every conjugacy class intersection with $D$ is a multiple of $w$, and the order of the group of linear characters is a power of $3$.
@@
Note that $4w^{2}-1$ is odd, so $G$ is a solvable group. Moreover, $\sqrt{\Delta}=2w$, $\theta_{1}=w$, and $\theta_{2}=-w$, so $v_{1}=2w-1$ and $v_{2}=2w+1$. Since $G$ is solvable, it has a nontrivial group of linear characters, so let $p$ be a prime dividing the order of this group. Then, by Theorem 3.14, either $v_{1}\equiv 1\pmod{p}$ and $v_{2}\equiv 0\pmod{p}$, or $v_{1}\equiv 0\pmod{p}$ and $v_{2}\equiv 1\pmod{p}$. After some algebraic manipulation, it is clear that $v_{1}\equiv 0\pmod{p}$ and $v_{2}\equiv 1\pmod{p}$ implies $p=1$, and that $v_{1}\equiv 1\pmod{p}$ and $v_{2}\equiv 0\pmod{p}$ implies $p=3$. Therefore, $3$ divides $2w+1$, implying $w\equiv 1\pmod{3}$. Additionally, by Theorem 3.14, the linear group of characters must be a power of $3$.
Now, let $a\in G$ be an element of order divisible by 33 such that $a\not\in[G,G]$. Such an element must exist, for 33 divides $|\mathcal{L}|$. In addition, let $\xi\in\mathcal{L}$. Then, by Theorem 3.12, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$. By Lemma 3.2, $\xi(D)=\pm w$. Set $w=3m+1$ for some $m\in\mathbb{Z}$. Therefore,
@@
Since $a$ has order divisible by $3$, then $3$ divides $|C_{G}(a)|$, so $6m+2-(\pm 1)\equiv 0\pmod{3}$. Since $w\neq 0\pmod{3}$, then $2-(\pm 1)\equiv 0\pmod{3}$. It follows that $\xi(D)=-w$, so that $2-(\pm 1)$ equals 33. ∎
@@ 4. Computational methods
These methods, combined with Theorem 3.12, makes it rather quick to search for groups which could contain a PDS just by utilizing the irreducible characters of the group. In some cases, simply knowing the values $|D\cap h^{G}|\pmod{p^{\ell}}$ can rule out the existence of a PDS for a certain parameter sets inside of a nonabelian group, such as in Example 4.2. In other cases, targeted information about $|D\cap h^{G}|$ can be used to quickly construct partial difference sets even inside of groups with up to $1000$ elements, using the search method described in [2].
@@ 4.1. Modular computation of conjugacy class intersection sizes for PDSs
Using the notation above, for $h\in G$ such that $h\neq 1$, suppose that $p^{\ell}$ does not divide $|C_{G}(h)|$. Then, over the ring $R$,
@@
There does not exist an invertible $(112,75,50)$-DS in a nonabelian group $G$ of order $112$. We suppose for contradiction that there does exist such a DS in a nonabelian group $G$ of order $112$. This parameter set has eigenvalues $\theta_{1}=5$ and $\theta_{2}=-5$. By Lemma 3.2, if $D$ exists then $\overline{\chi(D)}\equiv 0\pmod{5}$. Therefore, by Lemma 4.1, $\Phi(g)\equiv 75-5+\sum_{\chi\in{\rm Irr}(G)}(\chi(1)\theta_{2})\chi(g)\equiv 0\pmod{5}$. Finally, there is no centralizer $C_{G}(g)$ of order divisible by $5$, since $5$ does not divide $112$. Therefore, for all $g\in G$, $|g^{G}\cap D|\equiv 0\pmod{5}$.
Consulting GAP[15] on all possible nonabelian groups of order $112$, we may for a given group $G$ compute the largest multiple of $5$ less than or equal to the size of each conjugacy class of $G$. For example, if $G=$ SmallGroup(112, 41), then it has conjugacy class sizes $[1,1,8,7,8,7,8,8,8,8,8,8,8,8,8,8]$, and the largest multiple of 5 for each conjugacy class is $[0,0,5,5,5,5,5,5,5,5,5,5,5,5,5,5]$. In this case, $G$ cannot contain a reversible difference set, for Sum([0,0,5,5,5,5,5,5,5,5,5,5,5,5,5,5])=70<k=75[0,0,5,5,5,5,5,5,5,5,5,5,5,5,5,5])=70<k=75.
An analogous computation for each group verifies that there are no nonabelian groups of order $112$ containing a regular $(112,75,50)$-DS. A direct verification may be done in GAP[15], using the following command:
@@
Let $G$ contain a regular $(v,k,\lambda,\mu)$-PDS with eigenvalues $\theta_{1}>\theta_{2}$ such that $p^{\ell}$ divides both $\theta_{1}$ and $\theta_{2}$, then $k\equiv 0\pmod{p^{\ell}}$. In particular, $k\equiv 0\pmod{\gcd(\theta_{1},\theta_{2})}$.
@@
Let $G$ be a group containing a regular $(v,k,\lambda,\mu)$-PDS $D$ with eigenvalues $\theta_{1}>\theta_{2}$ such that $\gcd(\theta_{1},\theta_{2})\neq 1$. If $p^{\ell}$ divides $\gcd(\theta_{1},\theta_{2})$ and $p^{\ell}$ does not divide $v$, then for each conjugacy class $h^{G}$,
@@
We explore the structure of a $(183,112,66,72)$-PDS $D$ in a group $G$ of order $183$. The eigenvalues of this PDS are $\theta_{1}=4$ and $\theta_{2}=-10$. Then, $\gcd(4,-10)=2$, but $2$ does not divide $183$. Therefore, by Lemma 4.4, for every $h\in G$, we have $|h^{G}\cap D|\equiv 0\pmod{2}$. By a computation in GAP[15], there is one nonabelian group given by SmallGroup(183, 1) with two conjugacy classes $h_{1}^{G},h_{2}^{G}$ of order $61$, 2020 conjugacy classes of order $3$, and one conjugacy class of order $1$. If $x\in h_{1}^{G},h_{2}^{G}$, then $x\not\in G^{\prime}$ by a computation in GAP[15]. Moreover, every prime $p$ dividing $183$ does not divide $14$, and $G$ must have a subgroup of linear characters, for it is solvable by Feit-Thompson. So, by Theorem 3.12, $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|\in\{\frac{112-4}{3},\frac{112+10}{3}\}$. Only $\frac{112-4}{3}=36$ is integral, so $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|=36$. Therefore, 11272=40112-72=40 elements must be partitioned among $20$ conjugacy classes. The maximum size of each conjugacy class is $3$, and each class must have $0\pmod{2}$ elements, so each of the remaining conjugacy classes contain $2$ elements of $D$.
@@
Combining Lemma 4.1 and Theorem 3.12 lets us build a fast algorithm to solve for all values $|D\cap h^{G}|$ even in comparatively large groups of order up to $1000$. We briefly describe the concept of such an algorithm, and then provide the pseudocode for an implementation.
@@
We show that there does not exist a $(100,33,14,9)$-PDS inside of $G=$ SmallGroup(100, 3). By a computation in GAP[15], $G$ contains a conjugacy class $h^{G}$ of order $4$ such that $g,g^{-1}\in h^{G}$ for all $g\in h^{G}$. Moreover, the centralizer $C_{G}(h)$ has order $25$. Since the eigenvalues corresponding to $(100,33,14,9)$ are $8$ and 3-3, by Lemma 4.1,
@@
The only nonnegative integer less than or equal to $4$ that is congruent to $1\pmod{11}$ is $1$, so $|h^{G}\cap D|=1$. Then, by Lemma 4.6, $G$ cannot contain a $(100,33,14,9)$-PDS.